The following were obtained by applying Kirchoff’s laws to an electric circuit


2IA +IB −IC = −8


−IA +IB +IC = 3


−2IA +4IC = 18


1
Expert's answer
2022-06-16T09:49:51-0400

Augmented matrix


"\begin{pmatrix}
 2 & 1 & -1 & & -8 \\
 -1 & 1 & 1 & & 3 \\
 -2 & 0 & 4 & & 18 \\
\end{pmatrix}"

"R_1=R_1\/2"


"\begin{pmatrix}
 1 & 1/2 & -1/2 & & -4 \\
 -1 & 1 & 1 & & 3 \\
 -2 & 0 & 4 & & 18 \\
\end{pmatrix}"

"R_2=R_2+R_1"


"\begin{pmatrix}
 1 & 1/2 & -1/2 & & -4 \\
0 & 3/2 & 1/2 & & -1 \\
 -2 & 0 & 4 & & 18 \\
\end{pmatrix}"

"R_3=R_3+2R_1"


"\begin{pmatrix}
 1 & 1/2 & -1/2 & & -4 \\
0 & 3/2 & 1/2 & & -1 \\
 0 & 1 & 3 & & 10 \\
\end{pmatrix}"

"R_2=2R_2\/3"


"\begin{pmatrix}
 1 & 1/2 & -1/2 & & -4 \\
0 & 1 & 1/3 & & -2/3 \\
 0 & 1 & 3 & & 10 \\
\end{pmatrix}"

"R_1=R_1-R_2\/2"


"\begin{pmatrix}
 1 & 0 & -2/3 & & -11/3 \\
0 & 1 & 1/3 & & -2/3 \\
 0 & 1 & 3 & & 10 \\
\end{pmatrix}"

"R_3=R_3-R_2"


"\begin{pmatrix}
 1 & 0 & -2/3 & & -11/3 \\
0 & 1 & 1/3 & & -2/3 \\
 0 & 0 & 8/3 & & 32/3 \\
\end{pmatrix}"

"R_3=3R_3\/8"


"\begin{pmatrix}
 1 & 0 & -2/3 & & -11/3 \\
0 & 1 & 1/3 & & -2/3 \\
 0 & 0 & 1 & & 4 \\
\end{pmatrix}"

"R_1=R_1+2R_3\/3"


"\begin{pmatrix}
 1 & 0 & 0 & & -1 \\
0 & 1 & 1/3 & & -2/3 \\
 0 & 0 & 1 & & 4 \\
\end{pmatrix}"

"R_2=R_2-R_3\/3"


"\begin{pmatrix}
 1 & 0 & 0 & & -1 \\
0 & 1 & 0 & & -2\\
 0 & 0 & 1 & & 4 \\
\end{pmatrix}"

"I_A=-1, I_B=-2, I_C=4"